How do you rewrite a function in the form of f(x)=a(x-h)^2+k then complete the square of f(x)=2x^2-20x+3
Answer by nerdybill(7384)

How do you rewrite a function in the form of f(x)=a(x-h)^2+k then complete the square of f(x)=2x^2-20x+3
.
f(x)=2x^2-20x+3
group:
f(x)=(2x^2-20x)+3
factor group:
f(x)=2(x^2-10x)+3
take half 'b' square it and add:
f(x)=2(x^2-10x+25)+3-50
f(x)=2(x-5)(x-5)+3-50
f(x)=2(x-5)^2-47 (this is in the form of f(x)=a(x-h)^2+k)